Gary is currently 31 years older than his daughter. In 30 years' time he will be twice his daughter's age. Write an equation and use it to find the current ages of Gary and his daughter.
Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

let x be Gary's daughter's age
:
Gary's age is x+31
:
(x+31) +30 = 2 * (x+30)
:
x+61 = 2x +60
:
x = 1
:
***************************************************
Gary is 32 years old and his daughter is 1 year old
:
In 30 years Gary is 62 and his daughter is 31
:
answer checks
**************************************************
:

Answer by Alan3354(69443)

Gary is currently 31 years older than his daughter.
---------
"currently" ????
Is he not always 31 years older?
-----
Another instance of trying to add words to make someone seem smarter.
It doesn't work.